# Intumescent Polymer-Particle Hydrogels Expand the Retardancy Window for Wildfire Defense

### v1

@@ 2.4 Comparison to conventional intumescent char
It is useful to compare this novel PP hydrogel fire retardant with conventional intumescent char systems, as both enhance fire resistance through volume expansion and thermal insulation mechanisms. Conventional intumescent chars rely on a multi-component chemical reaction triggered at high temperatures, consisting of (i) an acid source or dehydrating agent, (ii) a carbon source or charring agent, and (iii) a blowing agent that generates gas to induce expansion.24, 31 While widely used in fire protection, these materials suffer from several critical limitations. One of the primary challenges is the significant shrinkage of the char layer under prolonged heat exposure. This shrinkage and cracking alter the pore size distribution, leading to increased thermal conductivity and reduced fire protection performance.32 Furthermore, non-continuous and brittle chars are mechanically fragile and prone to delamination, exposing the underlying substrate to direct flame contact and increasing the risk of failure in high-temperature environments. This degradation undermines the char’s structural integrity and its ability to provide effective thermal insulation.33
A major limitation of traditional intumescent coatings is the uncontrolled nature of their expansion, as the degree of foaming is highly dependent on the decomposition kinetics of the blowing agent34 and acid source.35 Variability in gas release and char formation rates can lead to inconsistent expansion, potentially compromising the coating’s protective performance. This inconsistency makes it difficult to achieve a uniform and mechanically stable intumescent layer, limiting the long-term fire protection effectiveness of these materials. These drawbacks highlight the need for more structurally robust fire-retardant materials, particularly for applications requiring extended fire resistance.
@@ 3 Conclusion
In this study, we demonstrated that a dynamically crosslinked hydrogel comprising robust PP interactions exhibits exceptional foaming ability and fire retardancy, even under dehydration, making it a promising alternative to conventional superabsorbant polymer based water-enhancing hydrogels. Unlike traditional formulations that rely primarily on water retention for fire suppression, the PP hydrogel system leverages an integrated self-healing and intumescent mechanism to achieve superior fire protection. This intumescent process is driven by the controlled degradation of cellulose ether, which releases gaseous byproducts that facilitate expansion, while concurrent silica densification stabilizes the aerogel structure, yielding a mesoporous, uniformly layered thermal barrier.
Notably, this approach replaces the need for conventional intumescent coatings that depend on a complex combination of charring agents, blowing agents, and dehydrating agents. Instead, the PP hydrogel offers a more streamlined, environmentally friendly, and scalable solution that achieves the same, or superior, thermal insulation and fire resistance through a facile composition. The formation of a stable aerogel layer upon heat exposure ensures long-lasting protection, mitigating structural damage in high-risk fire conditions.
@@ 4.4 Thermogravimetric Analysis
4.5 Burn Tests
A 20 g hydrogel sample was applied to a plywood plate, covering a 5 cm 5 cm square area (Lowe’s Unfinished Whitewood Board, cut into 10 cm 10 cm squares). A Bernzomatic MAP-Pro Premium Hand Torch, fueled by propylene (99.5-100 vol%) and propane (0-0.5 vol%), capable of reaching temperatures up to , was used for burning tests. The torch was stabilized approximately 5 cm above the hydrogel surface. Burn time was measured from the moment the torch was ignited to the point when the wood substrate visibly began to char.
@@ 4.8 Rheological Characterization
The authors appreciate the Gordon & Betty Moore Foundation for their financial support of this work as part of our efforts to develop wildland fire retardants to improve the execution of prescribed burns and enhance forest management (Grant Numbers 11539). Part of this work was performed at the Stanford Nano Shared Facilities (SNSF), supported by the National Science Foundation under award ECCS-2026822. Part of the rheology data was collected at the Stanford Soft and Hybrid Materials Facility (SMF), supported by the NSF grant National Science Foundation grant ECCS1542152.
@@
Data on the flow rheology of hydrogels, evaporation properties, thermal imaging of burn tests, and conductivity modeling are in the Supplemental Information. Burn tests are shown in the Supplementary Videos S1-S3.

### v2

@@ 2.4 Comparison to conventional intumescent char
It is useful to compare this novel PP hydrogel fire retardant with conventional intumescent char systems, as both enhance fire resistance through volume expansion and thermal insulation mechanisms. Conventional intumescent chars rely on a multi-component chemical reaction triggered at high temperatures, consisting of (i) an acid source or dehydrating agent, (ii) a carbon source or charring agent, and (iii) a blowing agent that generates gas to induce expansion.24, 31 While widely used in passive fire protection applications for buildings, furniture, and industrial infrastructure, these materials suffer from several critical limitations. One of the primary challenges is the significant shrinkage of the char layer under prolonged heat exposure. This shrinkage and cracking alter the pore size distribution, leading to increased thermal conductivity and reduced fire protection performance.32 Furthermore, non-continuous and brittle chars are mechanically fragile and prone to delamination, exposing the underlying substrate to direct flame contact and increasing the risk of failure in high-temperature environments. This degradation undermines the char’s structural integrity and its ability to provide effective thermal insulation.33 A major limitation of traditional intumescent coatings is the uncontrolled nature of their expansion, as the degree of foaming is highly dependent on the decomposition kinetics of the blowing agent34 and acid source.35 Variability in gas release and char formation rates can lead to inconsistent expansion, potentially compromising the coating’s protective performance. This inconsistency makes it difficult to achieve a uniform and mechanically stable intumescent layer, limiting the long-term fire protection effectiveness of these materials. These drawbacks highlight the need for more structurally robust fire-retardant materials, particularly for applications requiring extended fire resistance, such as wildfire protection and suppression.
@@ 3 Conclusion
In this study, we demonstrated that a dynamically crosslinked hydrogel comprising robust PP interactions exhibits exceptional foaming ability and fire retardancy, even under dehydration, making it a promising alternative to conventional superabsorbent polymer based water-enhancing hydrogels. Unlike traditional formulations that rely primarily on water retention for fire suppression, the PP hydrogel system leverages an integrated self-healing and intumescent mechanism to achieve superior fire protection. This intumescent process is driven by the controlled degradation of cellulose ether, which releases gaseous byproducts that facilitate expansion, while concurrent silica densification stabilizes the aerogel structure, yielding a mesoporous, uniformly layered thermal barrier.
Notably, these PP hydrogel materials are complementary to conventional intumescent coatings for structure protection. While traditional intumescent coatings may be applied to structures to harden them to wildfire, PP hydrogels function similar to other WEGs and are applied during wildfire scenarios to protect from oncoming fire. These PP hydrogels have exhibit the added benefit of a heat-activated hydrogel-to-aerogel transition that enhances the wildfire defense window, enabling longer-lasting protection against structural damage from wildfire.
@@ 4.4 Thermogravimetric Analysis
4.5 Wildfire-Relevant Flame Tests
We utilized small-scale burn tests adapted from ASTM E1321-1997(02), as recommended by the U.S. Forest Service’s Wildland Fire Chemicals program. 17, 37 A 20 g hydrogel sample was applied to a plywood plate, covering a 5 cm 5 cm square area (Lowe’s Unfinished Whitewood Board, cut into 10 cm 10 cm squares). A Bernzomatic MAP-Pro Premium Hand Torch, fueled by propylene (99.5-100 vol%) and propane (0-0.5 vol%), capable of reaching temperatures up to , was used for burning tests. The torch was stabilized approximately 5 cm above the hydrogel surface. Burn time was measured from the moment the torch was ignited to the point when the wood substrate visibly began to char.
@@ 4.8 Rheological Characterization
The authors appreciate the Gordon & Betty Moore Foundation for their financial support of this work as part of our efforts to develop wildland fire retardants to improve the execution of prescribed burns and enhance forest management (Grant Numbers 11539). Part of this work was performed at the Stanford Nano Shared Facilities (SNSF), supported by the National Science Foundation under award ECCS-2026822. Part of the rheology data was collected at the Stanford Soft and Hybrid Materials Facility (SMF), supported by the NSF grant National Science Foundation grant ECCS1542152. The authors gratefully acknowledge the support of the Stanford Strategic Energy Research Consortium (Award No. 327778) for Brunauer–Emmett–Teller (BET) surface area measurements.
@@
Data on the flow rheology of hydrogels, evaporation properties, thermal imaging of burn tests, and conductivity modeling are in the Supplemental Information. Burn tests are shown in the Supplementary Videos S1-S3. Video S4 shows the difference in flow properties of AquaGK and MHEC/CSP hydrogels, with water as the control.